You propose a new temperature scale with temperatures given in ∘M. You define 0.0 ∘M to be the normal melting point of mercury and 100.0 ∘M to be the normal boiling point of mercury.

A. What is the normal boiling point of water in ∘M?

B. A temperature change of 16.0 ∘M corresponds to how many ∘C?
Answer #1

Delta Tc/delta Tm=(boiling point in C-freezing point inC)/(boiling point in M- freezing point in M)

Del.Tc/Del.Tm=(357-(-39))/(100-0)=del.Tc=del.Tm×3.96

Del.Tc=63.36°M corresponding to 16°M

Tm=(Tc+39℃)100/396=139×100/396=35.101°M